### v1

# SPGCL: Effective Graph Contrastive Learning via SVD-Guided Structural Perturbation

Abstract
Graph Neural Networks (GNNs) can be highly sensitive to structural noise, including spurious or missing edges caused by adversarial attacks or non-adversarial imperfections. Existing graph contrastive learning methods rely on either random perturbation (e.g., edge dropping) to generate diverse views or purely spectral augmentation (e.g., SVD) to preserve global structural priors. However, random perturbations are structure-agnostic and risk removing critical edges, while purely SVD-based views tend to be dense and lack sufficient diversity. To bridge this gap, we propose SPGCL - a robust graph contrastive learning framework via SVD-guided structural perturbation. Our method couples a lightweight stochastic edge removal with an SVD-guided refinement. This design not only retains the diversity of random augmentations and overcomes the densification issue of spectral methods but also delivers three key advantages: (1) Robustness to accidental deletion: important edges mistakenly removed in the first stage are likely to be recovered by the subsequent SVD-guided step, mitigating harmful effects while preserving diversity. (2) Enrichment with missing links: the SVD-based scoring can introduce new, semantically meaningful edges absent in the original graph, making the contrastive view more informative. (3) Controllable structural discrepancy: by balancing the rates of edge removal and recovery, we explicitly control the edge count between views, ensuring that the contrastive signal stems from semantic structural differences rather than mere edge-number gaps. To further align the two views, we incorporate a contrastive fusion module regularized by a global similarity constraint. Extensive experiments on ten benchmark datasets demonstrate that SPGCL consistently improves the robustness and accuracy of base GNNs, outperforming state-of-the-art graph contrastive learning and structure learning methods.
I Introduction
Graph Neural Networks (GNNs) have become the de facto standard for learning representations from relational data, achieving remarkable success in domains ranging from social network analysis to biochemical discovery [10, 24, 8, 21, 27]. Their core strength lies in recursively aggregating information from neighboring nodes, effectively leveraging both node attributes and graph topology.
However, this strength is also a critical vulnerability: GNNs are highly sensitive to imperfections in the input graph structure [35]. In practice, graphs are often noisy due to adversarial attacks [35, 3, 25] (e.g., fake links in recommendation systems) [29, 22] or non-adversarial imperfections (e.g., missing observations or measurement errors). Such structural noise can corrupt the message-passing process, leading to severe degradation in model performance and robustness.
To mitigate this issue, Graph Contrastive Learning (GCL) has emerged as a powerful self-supervised paradigm. By contrasting different “views” of the same graph, GCL encourages the model to learn representations that are invariant to benign perturbations [14, 30]. The crux of GCL lies in designing effective graph augmentation strategies to generate these views. Prevailing methods largely fall into two categories. The first employs random perturbations, such as uniformly dropping edges [33]. While simple and conducive to view diversity, these operations are structure-agnostic and may inadvertently remove semantically critical edges, providing noisy or even detrimental contrastive signals. The second category leverages spectral methods, most notably using Singular Value Decomposition (SVD) to generate a globally informed view [2]. This approach preserves high-order structural patterns but often results in a dense and relatively deterministic view, which may limit the diversity crucial for contrastive learning and introduce many weak connections.
This dichotomy presents a clear gap: we need a perturbation strategy that is both structure-aware and diversity-preserving. To bridge this gap, we propose SPGCL, a novel framework for robust graph contrastive learning via SVD-guided structural perturbation. Our core innovation is a coupled two-stage augmentation strategy: a lightweight stochastic edge removal followed by an SVD-guided scoring and sparse edge recovery. This design is not merely a combination but a synergistic integration that delivers distinct advantages:
Risk Mitigation for Random Deletion: If the initial random drop accidentally removes an important edge, our SVD-guided recovery step can, with high probability, restore it. This preserves the beneficial diversity introduced by randomness while dramatically reducing its potential harm [12].
Enrichment Beyond the Original Graph: Beyond recovery, the SVD-based scoring can identify and add new, semantically meaningful edges that are missing from the original graph, thereby creating a more informative contrastive view.
Controlled Structural Discrepancy: By balancing the rates of edge removal and recovery, we can explicitly control the total edge count in the perturbed view. This ensures the contrastive signal stems from meaningful structural semantics rather than trivial differences in edge density.
Furthermore, we introduce a contrastive fusion module with a global similarity-matrix constraint to align node embeddings from the original and perturbed views, ensuring consistency in both local and global structural patterns.
Extensive experiments on ten benchmark datasets demonstrate that SPGCL consistently outperforms state-of-the-art GCL and graph structure learning baselines. Ablation studies confirm the contribution of each component, and robustness analyses show our model’s superior stability under increasing structural noise.
Our main contributions are summarized as follows:
We propose a novel SVD-guided structural perturbation strategy for GCL that synergistically couples stochastic edge removal with spectral refinement, addressing the limitations of both random and purely spectral augmentations.
Our design delivers three key benefits: robustness to accidental edge deletion, enrichment via missing link addition, and controlled structural discrepancy, advancing the design of graph augmentations.
We incorporate a contrastive fusion scheme with a global consistency constraint to effectively align multi-view representations.
We provide comprehensive empirical evidence of SPGCL’s effectiveness and robustness across diverse real-world graphs.
II Related Work
II-A Robustness in Graph Neural Networks
The sensitivity of GNNs to structural noise has spurred extensive research into robust learning. Early adversarial defense works focus on identifying and hardening models against malicious perturbations [35, 25, 5, 19, 26, 32]. Another line of work, Graph Structure Learning (GSL), aims to infer an optimal or denoised graph topology jointly with GNN training. Methods like LDS [7] and Pro-GNN [9] impose constraints (e.g., low-rank, sparsity) to reconstruct cleaner graphs. While improving robustness, these methods often operate in a single-view manner and may not explicitly preserve the diverse, task-agnostic signals crucial for generalizable representation learning.
II-B Graph Contrastive Learning and Augmentation Strategies
Graph Contrastive Learning (GCL) provides a self-supervised pathway to robust representations by contrasting multiple augmented views of the graph [30, 33]. The performance of GCL hinges critically on the design of graph augmentation. Existing augmentation strategies can be categorized along a spectrum from local/random to global/spectral:
Local Random Perturbations: Most GCL frameworks [33, 34] rely on random operations like edge dropping and feature masking. These are simple and promote view diversity. However, their structure-agnostic nature is a fundamental limitation: they risk removing important edges and fail to inject semantically meaningful structural information, making the contrastive signal potentially noisy or uninformative.
Global Spectral Augmentations: To capture higher-order and global structural patterns, recent methods like LightGCL [2] employ Singular Value Decomposition (SVD) to generate augmented views. While more structure-aware, purely SVD-based views tend to become dense and deterministic, which can reduce the effective diversity between views and introduce many weak connections.
This dichotomy highlights a critical gap: existing methods either prioritize diversity at the cost of structural rationality (random) or prioritize global consistency at the cost of diversity and sparsity (spectral). A principled integration that harnesses the benefits of both paradigms remains underexplored.
II-C Beyond Diversity: Towards Rational and Controlled Perturbation
Recent studies have begun to scrutinize the quality of augmentations. Some works point out the instability of random augmentations [31], while others attempt to guide perturbations with node features or gradients [34]. However, these approaches often lack a global structural prior or do not address the post-perturbation recovery of critical edges. Crucially, few methods consider explicitly controlling the structural discrepancy between views (e.g., edge count), leaving open the risk that contrastive learning might exploit trivial, task-irrelevant differences. Our work directly addresses these shortcomings. SPGCL introduces a novel SVD-guided structural perturbation that not only couples random diversity with spectral rationality but also provides a recovery mechanism for accidentally deleted edges, enriches the view with missing high-order links, and enables explicit control over the perturbation scale. This positions SPGCL as a step towards more principled, interpretable, and effective graph augmentation for contrastive learning.
III SVD-Guided Structural Perturbation
This section details the core component of SPGCL: an SVD-guided structural perturbation strategy designed to generate an augmented view that is both diverse and structure-aware. We first introduce the necessary preliminaries, then elaborate on the motivation and the step-by-step perturbation process.
III-A Preliminaries
III-A1 Problem Setting
We consider an attributed graph , where is the node set with , is the edge set, and is the node feature matrix. The graph structure is represented by the adjacency matrix , where if and otherwise. We focus on semi-supervised node classification, where only a subset of nodes are labeled.
III-A2 Graph Convolutional Network Backbone
We adopt the Graph Convolutional Network (GCN) [10] as our encoder. Each layer propagates information as:
where , is the diagonal degree matrix of , and is a nonlinear activation.
III-A3 Low-Rank Approximation via SVD
Singular Value Decomposition (SVD) provides a low-rank approximation of the adjacency matrix:
where are orthogonal matrices, with . A rank- truncated SVD retains the top- components to capture the most significant global structural patterns:
III-B Motivation and Design Intuition
Existing augmentation strategies for graph contrastive learning suffer from a fundamental trade-off: random perturbations (e.g., edge dropping) are diverse but structure-agnostic, while purely spectral augmentations (e.g., SVD) are structure-aware but tend to be dense and deterministic. Our goal is to generate a perturbed view that simultaneously achieves three objectives:
1.
Preserving diversity for effective contrastive learning,
2.
Maintaining global structural consistency by avoiding arbitrary corruption of important edges,
3.
Ensuring sparsity to avoid introducing excessive weak connections.
To this end, we propose a two-stage structured perturbation that couples stochastic edge removal with SVD-guided refinement. This design not only balances diversity and structural awareness but also introduces three key advantages:
Robust recovery of accidentally removed edges: Important edges dropped in the first stage can be recovered via SVD-guided scoring.
Enrichment with missing high-order connections: The SVD-based scoring can identify and add semantically meaningful edges absent in the original graph.
Controllable structural discrepancy: By balancing removal and addition, we explicitly control the edge count, ensuring the contrastive signal stems from semantic differences rather than mere edge-number gaps.
III-C Two-Stage Structured Perturbation
Our perturbation process, illustrated in Fig. 1, consists of two synergistic stages.
Stage 1: Lightweight Stochastic Edge Removal.
We first randomly remove a small fraction of edges from to form a removed set . The remaining edge set is , with corresponding adjacency matrices and , such that . This step injects necessary diversity into the view.
Stage 2: SVD-Guided Scoring and Sparse Recovery.
We then perform truncated SVD on :
To account for the structural impact of the removed edges, we compute a first-order approximation of the singular-value shift ([11, 4]):
A refined score matrix is constructed by adjusting the singular values:
This matrix encodes global structural information while reflecting the perturbation.
Top- sparsification. Importantly, we sparsify to avoid densification [2]. We rank candidate edges not in by their scores in , select the top- entries, and form a sparse addition matrix . The number of recovered edges is determined by:
where is the edge drop ratio, and it also serves as the base scale for edge recovery and is a flexible adjustment parameter [4]. In practice, can be positive, zero, or negative, thereby controlling whether we recover more, the same, or fewer edges than those initially removed.
The final perturbed adjacency is:
where controls the strength of added edges. This yields a sparse, structure-aware augmented view ready for contrastive learning.
Summary of Advantages.
Our two-stage perturbation directly addresses the aforementioned objectives:
The random removal ensures diversity.
The SVD-guided scoring preserves global structural patterns and enables recovery of important edges.
The top- sparsification maintains sparsity and allows explicit control over the edge count.
This principled design forms the foundation of SPGCL, enabling robust and informative contrastive learning even under structural noise.
IV The Proposed Method: SPGCL
In this section, we present the overall architecture of SPGCL, which consists of two key components: (1) SVD-guided structural perturbation for generating a sparse, structure-aware augmented view, and (2) contrastive learning-based fusion for aligning the original and perturbed views while enforcing global structural consistency. The complete framework is illustrated in Fig. 1.
IV-A Framework Overview
SPGCL follows a dual-view contrastive learning paradigm. Given an input graph with adjacency matrix , we first generate a perturbed view via the structured perturbation described in Section III. Both the original graph (with adjacency ) and the perturbed graph (with adjacency ) are then encoded by a shared GNN encoder to obtain node embeddings. Finally, a contrastive fusion module aligns the two views at both node and global structural levels.
IV-B Contrastive Learning-Based Fusion
While the SVD-guided perturbation generates a structurally informed view, the inherent discrepancy between the original and perturbed views may lead to inconsistent node representations if not properly aligned. To address this, we introduce a contrastive fusion scheme that operates at two levels: node-level alignment via InfoNCE loss and global structural alignment via a consistency constraint.
Node-Level Alignment via InfoNCE
We employ a Siamese GNN encoder to extract layer-wise node embeddings from both views. Let and denote the embeddings of node at the -th layer from the perturbed and original views, respectively. Following [14], we apply the InfoNCE loss at each layer to maximize the agreement between corresponding node embeddings across views:
where denotes cosine similarity and is a temperature hyperparameter. This loss encourages the model to learn node representations that are invariant to the applied structural perturbations.
Global Structural Alignment via Consistency Constraint
The InfoNCE loss aligns individual node embeddings but does not explicitly preserve the global similarity structure among nodes. To ensure that the perturbed view retains the overall topological patterns of the original graph, we introduce an additional consistency constraint [23]. Specifically, we first normalize the final-layer embeddings (-th layer) of both views:
where
and denotes the -norm of vectors. We then compute the node similarity matrices for both views as:
These matrices encode the global pairwise similarity structure within each view. To enforce consistency, we minimize the Frobenius norm of their difference:
This constraint ensures that the perturbed view does not deviate arbitrarily from the original graph’s global relational pattern, thereby preserving structural integrity.
Joint Optimization Objective
The overall loss function for SPGCL combines the node-level contrastive loss and the global consistency constraint:
where is the supervised classification loss (e.g., cross-entropy) on the labeled nodes, and are hyperparameters balancing the contributions of the contrastive and consistency terms. During training, all components are optimized jointly, enabling the model to learn robust node representations that are both discriminative for the downstream task and invariant to structure-preserving perturbations.
V Experiments
V-A Experimental Setup
V-A1 Datasets
We evaluate SPGCL and baseline models on ten real-world datasets spanning various domains and structural properties. The statistics are summarized in Table I. These include:
Social networks (BlogCatalog, Flickr) [28]: Nodes are users and edges indicate friendships or followships.
Co-purchase network (Amazon-Photo) [18]: Nodes are products and edges indicate frequent co-purchases.
Actor co-occurrence network (Actor) [15]: Nodes are actors and edges denote co-appearance in films.
The datasets cover a wide range of sizes, densities, and homophily levels, providing a comprehensive testbed for evaluating robustness and generalization.
| Dataset | Nodes() | Edges() | Classes | Features() | Train nodes | Val/Test nodes |
| Cora | 2,708 | 5,429 | 7 | 1,433 | 140 | 500/1,000 |
| Citeseer | 3,327 | 4,732 | 6 | 3,703 | 120 | 500/1,000 |
| CoraML | 2,995 | 16,316 | 7 | 2,879 | 1,797 | 599/599 |
| Chameleon | 2,277 | 36,101 | 5 | 2,325 | 1,092 | 729/456 |
| Squirrel | 5,201 | 217,073 | 5 | 2,089 | 2,496 | 1,644/1,041 |
| Wiki | 2,405 | 17,981 | 17 | 4,973 | 1,443 | 481/481 |
| BlogCatalog | 5,196 | 343,486 | 6 | 8,189 | 3,117 | 1,039/1,040 |
| Flickr | 7,575 | 479,476 | 9 | 12,047 | 4,545 | 1,515/1,515 |
| Amazon-Photo | 7,650 | 238,162 | 8 | 745 | 4,590 | 1,530/1,530 |
| Actor | 7,600 | 30,019 | 5 | 932 | 3,648 | 2,432/1,520 |
V-A2 Baselines
V-A3 Implementation Details
All experiments are conducted on a Linux server with an NVIDIA GeForce RTX 3090 GPU and an AMD EPYC 7282 CPU. We implement SPGCL in Python 3.12.9 using PyTorch 2.5.1. For fairness, we use the same GCN backbone for all GNN-based methods and follow the same semi-supervised splitting strategy as in [6]. Hyperparameters are tuned via validation set performance. We report the average accuracy over independent runs.
V-B Node Classification Performance
V-B1 Overall Comparison
Table II presents the node classification accuracy across ten datasets. We observe:
Traditional GNNs (GCN, SGC, GraphSAGE, GAT) perform well on homophilous graphs (Cora, Citeseer) but degrade significantly on heterophilous ones (Chameleon, Squirrel, Actor).
Graph structure learning methods show improved robustness on some heterophilous datasets (e.g., Pro-GNN on Chameleon), but their performance is inconsistent across graphs.
DropEdge helps alleviate over-smoothing and achieves competitive results on some social networks (BlogCatalog, Flickr) but struggles with complex heterophily.
SPGCL achieves the best or second-best performance on nearly all datasets. It exhibits particularly large gains on challenging heterophilous graphs (e.g., over Pro-GNN on Chameleon, on Squirrel) while maintaining strong accuracy on homophilous ones.
These results validate that our structured perturbation effectively preserves important structural patterns while introducing beneficial diversity, leading to more robust representations.
| Method | Cora | Citeseer | CoraML | Chameleon | Squirrel | Wiki | BlogCatalog | Flickr | Amazon-Photo | Actor |
| GCN | 80.10 | 70.20 | 80.13 | 35.21 | 27.74 | 53.66 | 62.75 | 59.54 | 55.23 | 28.91 |
| SGC | 80.00 | 71.40 | 32.05 | 34.64 | 27.87 | 61.12 | 42.88 | 16.37 | 25.36 | 26.66 |
| GraphSAGE | 80.90 | 71.50 | 89.98 | 47.34 | 33.74 | 59.88 | 69.42 | 60.94 | 87.58 | 33.84 |
| GAT | 80.80 | 68.90 | 86.14 | 44.12 | 27.28 | 55.30 | 57.88 | 46.83 | 84.97 | 28.54 |
| LDS | 80.60 | 71.12 | 71.62 | 36.75 | 27.87 | 45.32 | 77.79 | 58.48 | 52.61 | 28.56 |
| Pro-GNN | 81.30 | 71.50 | 85.48 | 56.20 | 34.97 | 74.84 | 78.08 | 57.42 | 70.59 | 30.83 |
| Geom-GCN | 70.80 | 59.17 | 76.79 | 26.07 | 24.37 | 36.01 | 47.79 | 37.69 | 25.32 | 25.31 |
| DropEdge | 80.70 | 70.90 | 85.81 | 32.42 | 25.27 | 77.34 | 75.96 | 61.58 | 75.40 | 28.62 |
| Ours | 83.10 | 74.30 | 90.32 | 61.84 | 38.62 | 78.69 | 79.62 | 58.04 | 90.84 | 30.92 |
V-B2 Comparison with Different Perturbation Strategies
To isolate the benefit of our structured perturbation, we compare SPGCL against two naive perturbation strategies:
Node perturbation: Randomly perturb node features by adding Gaussian noise.
Edge perturbation: Randomly add or drop edges uniformly.
Results in Table V show that SPGCL significantly outperforms both naive strategies across all datasets, with an average improvement of . Edge perturbation generally performs better than node perturbation, confirming that structural noise is a more critical challenge in graph representation learning. The superior performance of SPGCL highlights the importance of coupling stochastic removal with SVD-guided recovery, which prevents arbitrary corruption and enriches the view with high-order connections.
V-C Ablation Study
Ablation on the perturbation module.
We conduct an ablation study to validate the design of our perturbation module, comparing SPGCL with three variants:
SVD: Use only the SVD-based reconstruction without stochastic removal.
Edge: Use only random edge dropping without SVD recovery.
SVD & Edge: Apply SVD first, then perform random edge dropping on the reconstructed view.
Results in Table III lead to several insights:
SPGCL outperforms all variants, demonstrating the necessity of coupling stochastic and spectral perturbations.
Edge alone often degrades performance, especially on heterophilous graphs, confirming that structure-agnostic random dropping is harmful.
SVD alone performs reasonably well, particularly on datasets like CoraML and Wiki, showing the value of global structural awareness.
SVD & Edge performs poorly across all datasets. This is because dropping edges after SVD reconstruction randomly destroys the globally consistent patterns recovered by SVD, resulting in a view that is neither diverse nor structure-aware.
These findings underscore that our two-stage design-where stochastic removal provides diversity and SVD-guided recovery preserves and enhances structural integrity-is crucial for generating effective contrastive views.
Ablation on the global consistency loss .
To evaluate the effect of the proposed global similarity-matrix consistency regularizer, we remove from the objective while keeping all other components unchanged. As shown in Table IV, discarding consistently degrades performance on both homophilous (Cora, Citeseer) and heterophilous datasets (Chameleon, Squirrel). This indicates that provides complementary supervision beyond the node-level InfoNCE alignment by stabilizing the global relational structure across the two views.
| Method | SVD | Edge | SVD&Edge | Ours |
| Cora | 81.30 | 75.40 | 71.00 | 83.10 |
| Citeseer | 67.00 | 60.30 | 57.60 | 74.30 |
| CoraML | 89.51 | 85.81 | 71.62 | 90.32 |
| Chameleon | 60.73 | 44.75 | 49.91 | 61.84 |
| Squirrel | 32.47 | 28.81 | 25.79 | 38.62 |
| Wiki | 72.77 | 57.08 | 60.08 | 78.69 |
| BlogCatalog | 61.54 | 66.80 | 38.75 | 79.62 |
| Flickr | 42.11 | 32.00 | 20.40 | 58.04 |
| Amazon-Photo | 88.23 | 82.29 | 47.52 | 90.84 |
| Actor | 24.94 | 24.75 | 22.41 | 30.92 |
| Variant | Cora | Citeseer | Chameleon | Squirrel |
| 83.10 | 74.30 | 61.84 | 38.62 | |
| 82.20 | 72.40 | 61.62 | 37.66 |
| Method | Node | Edge | Ours |
| Cora | 68.10 | 75.40 | 83.10 |
| Citeseer | 64.70 | 60.30 | 74.30 |
| CoraML | 85.00 | 85.81 | 90.32 |
| Chameleon | 32.12 | 44.75 | 61.84 |
| Squirrel | 25.47 | 28.81 | 38.62 |
| Wiki | 64.97 | 57.08 | 78.69 |
| BlogCatalog | 62.25 | 66.80 | 79.62 |
| Flickr | 32.35 | 32.00 | 58.04 |
| Amazon-Photo | 82.21 | 82.29 | 90.84 |
| Actor | 24.32 | 24.75 | 30.92 |
V-D Robustness Analysis
To evaluate robustness under structural noise, we simulate increasingly corrupted graphs by randomly removing and of edges, and compare the performance of SPGCL with baselines. Results in Table VI show:
Most baselines (GCN, SGC, GraphSAGE, GAT) suffer noticeable accuracy drops as the perturbation ratio increases.
SPGCL maintains the highest or competitive accuracy under both corruption levels, with a smaller relative decline. For example, on Cora, SPGCL drops by only when of edges are removed, whereas GCN drops by .
This resilience stems from our SVD-guided recovery, which can reconstruct important edges even under significant random deletion, and the contrastive fusion, which aligns representations across noisy views.
| Method | Cora | CiteSeer | CoraML | |||
| 0.2 | 0.3 | 0.2 | 0.3 | 0.2 | 0.3 | |
| GCN | 76.13 | 72.30 | 64.30 | 65.80 | 77.30 | 73.62 |
| SGC | 76.10 | 72.70 | 67.80 | 65.90 | 81.97 | 78.63 |
| SAGE | 74.80 | 71.70 | 67.10 | 64.00 | 85.98 | 82.47 |
| GAT | 74.80 | 71.70 | 66.30 | 63.50 | 79.97 | 73.62 |
| LDS | 74.90 | 72.00 | 67.30 | 64.10 | 59.43 | 52.25 |
| Gemo | 66.60 | 68.90 | 49.60 | 58.10 | 74.79 | 74.63 |
| DropEdge | 75.70 | 72.50 | 65.90 | 63.30 | 76.13 | 72.45 |
| Our | 80.50 | 74.14 | 66.20 | 64.40 | 88.98 | 76.74 |
V-E Visualization of Learned Embeddings
We visualize the test-set embeddings learned by different models on CoraML using t-SNE [20]. As shown in Fig. 2:
GCN, LDS, Geom-GCN, and DropEdge produce embeddings with blurred class boundaries and overlapping clusters.
SPGCL generates well-separated, compact clusters with clearer margins, indicating that the learned representations are more discriminative and structurally coherent.
The visualization confirms that our method better preserves intra-class similarity and inter-class distinction, aligning with the quantitative performance gains.
V-F Analysis of and
We study two hyper-parameters in SPGCL: the base edge generation ratio and an offset in Eq (7).
Setup.
We conduct a 2D grid search over
and report node classification accuracy for each pair. The corresponding heatmaps for Cora and Citeseer are shown in Fig. 3.
Observations.
We observe a clear hill-like (unimodal) trend over the grid. When Top- is too small, the perturbed view is overly similar to the original graph, resulting in limited view diversity. As increases to a moderate range, performance improves and reaches a peak region. However, overly large Top- introduces many low-quality edges, which may cause graph densification and degrade accuracy.
V-G Summary of Experimental Findings
Through extensive experiments, we demonstrate that:
1.
SPGCL consistently outperforms state-of-the-art GNNs and graph structure learning methods across diverse graph types.
2.
Our structured perturbation is superior to naive random perturbations, validating its design principles.
3.
Ablation studies confirm the importance of both stochastic removal and SVD-guided recovery.
4.
Ablating the global consistency loss leads to consistent performance drops on both homophilous and heterophilous datasets (Table IV), verifying that provides complementary global structural alignment beyond node-level contrastive learning.
5.
SPGCL exhibits stronger robustness under increasing structural noise.
6.
Visualization reveals that SPGCL learns more separable and semantically meaningful embeddings.
7.
Sensitivity analysis of shows a clear peak region (Fig. 3), indicating that moderate perturbation is optimal while under- or over-perturbation degrades performance.
These results collectively affirm the effectiveness of SPGCL in learning robust graph representations under structural noise and heterogeneity.
VI Conclusion
In this paper, we addressed the vulnerability of Graph Neural Networks (GNNs) to structural noise by proposing SPGCL, a robust graph contrastive learning framework via SVD-guided structural perturbation. Unlike random perturbations that are structure-agnostic or purely spectral augmentations that may become dense, SPGCL introduces a principled two-stage strategy: it first applies lightweight stochastic edge removal to ensure view diversity, then employs truncated SVD to recover and enhance topologically important edges. This design not only mitigates the risk of accidentally removing critical connections but also enriches the contrastive view with semantically meaningful edges missing in the original graph, while allowing explicit control over structural discrepancy.
Furthermore, we introduced a contrastive fusion module with a global consistency constraint, which aligns node embeddings across views at both local and global structural levels, ensuring robust and invariant representations. Extensive experiments on ten benchmark datasets demonstrate that SPGCL consistently outperforms state-of-the-art GNNs and graph structure learning methods, particularly on heterophilous graphs where structural noise is more detrimental. Ablation studies confirm the necessity of coupling stochastic and spectral perturbations, and robustness analysis shows SPGCL’s superior resilience under increasing structural corruption.
In summary, SPGCL provides an effective, generalizable framework for learning robust graph representations in noisy environments. Future work may explore adaptive perturbation ratios and extend the paradigm to dynamic or heterogeneous graphs.
VII Biography Section
If you have an EPS/PDF photo (graphicx package needed), extra braces are needed around the contents of the optional argument to biography to prevent the LaTeX parser from getting confused when it sees the complicated includegraphics command within an optional argument. (You can create your own custom macro containing the includegraphics command to make things simpler here.)